OpenStudy (mathmath333):

There are 20 people among whom two are sisters. Find the number of ways in which we can arrange them around a circle so that there is exactly one person between the two sisters?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ There are 20 people among whom two are sisters. }\hspace{.33em}\\~\\ & \normalsize \text{ Find the number of ways in which we can arrange }\hspace{.33em}\\~\\ & \normalsize \text{ them around a circle so that there is exactly one}\hspace{.33em}\\~\\ & \normalsize \text{ person between the two sisters?}\hspace{.33em}\\~\\ \end{align}}\)

jimthompson5910 (jim_thompson5910):

Think of 20 slots laid out like this  I can't fit all 20, but you get the idea

jimthompson5910 (jim_thompson5910):

the first and third slots are locked up by the 2 sisters 

jimthompson5910 (jim_thompson5910):

we have 20-2 = 18 slots left with 18 people left to fill them so there are 18*17*16*...*3*2*1 = 18! ways to fill up the rest of the slots now because there are 2 ways for the sisters to sit in the first and third seats, we double that to say 2*18! total ways to arrange the people

OpenStudy (mathmath333):

is answer 2*18!

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (mathmath333):

ok thanks